# HopaDIFF: Holistic-Partial Aware Fourier Conditioned Diffusion for Referring Human Action Segmentation in Multi-Person Scenarios

### v1

@@ Abstract
Action segmentation is a core challenge in high-level video understanding, aiming to partition untrimmed videos into segments and assign each a label from a predefined action set. Existing methods primarily address single-person activities with fixed action sequences, overlooking multi-person scenarios. In this work, we pioneer textual reference-guided human action segmentation in multi-person settings, where a textual description specifies the target person for segmentation. We introduce the first dataset for Referring Human Action Segmentation, i.e., RHAS133, built from $133$ movies and annotated with $137$ fine-grained actions with $33h$ video data, together with textual descriptions for this new task. Benchmarking existing action recognition methods on RHAS133 using VLM-based feature extractors reveals limited performance and poor aggregation of visual cues for the target person. To address this, we propose a holistic-partial aware Fourier-conditioned diffusion framework, i.e., HopaDIFF, leveraging a novel cross-input gate attentional xLSTM to enhance holistic-partial long-range reasoning and a novel Fourier condition to introduce more fine-grained control to improve the action segmentation generation. HopaDIFF achieves state-of-the-art results on RHAS133 in diverse evaluation settings. The code is available at https://github.com/KPeng9510/HopaDIFF.git.
@@ 1 Introduction
Human action recognition [6, 39, 35, 53, 42, 17] is a core problem with applications in robotics [41], surveillance [50], and healthcare [40]. While recent methods have shown strong performance, they typically rely on pre-trimmed video clips, which demand extensive manual pre-processing and limit applicability in real-time or continuous scenarios. Human action segmentation [13, 28, 54, 57], by contrast, operates on untrimmed video streams and aims to identify both action classes and their temporal boundaries. This setting is more challenging but aligns more closely with real-world use cases, where continuous and online interpretation of human behavior is essential.
@@
However, existing human action segmentation approaches [37, 31, 2, 15] are predominantly designed for single-person scenarios, where the performed actions follow a predefined sequence dictated by a specific task protocol, e.g., assembling a tool [44] based on fixed instructions or preparing a salad [47] according to a set menu. Such settings lack the complexity and variability present in real-world daily life scenarios, where multiple individuals may appear in the same untrimmed video, and the action sequences are not pre-defined but exhibit significant randomness and spontaneity. This limitation hinders the generalizability of current human action segmentation methods to more dynamic and unconstrained environments.
@@
In this work, we introduce a novel task, termed Referring Human Action Segmentation (RHAS), which, to the best of our knowledge, is the first to address action segmentation in multi-person untrimmed videos guided by textual references. We contribute the RHAS133 dataset, which comprises $133$ untrimmed videos featuring $542$ individuals, where each person of interest is annotated with a textual reference inspired by prior works on referring video and image understanding [39]. These textual descriptions serve as guidance for the model to localize and segment the actions of the referred individuals, with fine-grained, multi-label action annotations provided at the frame level following the AVA dataset format [18], as shown in Fig. 1.
@@
We further adapt several well-established human action segmentation methods, i.e., FACT [37], ActDiff [31], ASQuery [15], and LTContent [2], as well as one referring human action recognition method, i.e., RefAtomNet [39], to the proposed RHAS setting and evaluate their performance on our newly constructed dataset, which serves as the first RHAS benchmark. However, the experimental results reveal that these baseline methods achieve limited performance under the RHAS framework. This can be attributed to the fact that existing human action segmentation approaches are primarily designed for single-person scenarios without textual guidance and are not equipped to handle unstructured action sequences without a fixed protocol in complex, multi-person movie environments. Meanwhile, RefAtomNet [42], originally developed for referring action recognition in trimmed videos, lacks the capacity for temporal modeling and long video understanding required by the RHAS task.
To improve referring human action segmentation, we introduce HopaDIFF, a novel framework that employs a holistic- and partial-aware Fourier conditioned diffusion strategy guided by untrimmed video and textual references. It consists of two branches: a holistic branch that captures global context from the full video, and a partial branch that leverages GroundingDINO [34] to detect the referred person, generating a cropped video stream for fine-grained representation using a Vision-Language Model (VLM), e.g., BLIP-2 [26].
@@
To facilitate effective long video reasoning and holistic-partial aware information exchange, we introduce a novel module, i.e., HP-xLSTM, which fuses holistic and partial features using cross-input gate attentional xLSTM. Furthermore, to exert finer control over the generative process of the diffusion model for action segmentation, we incorporate an additional conditioning mechanism from the frequency domain based on the embeddings delivered by the HP-xLSTM. This is achieved by applying a Discrete Fourier Transform (DFT) along the temporal axis to extract frequency-based cues, which are combined with original feature conditions to guide the denoising process. The final segmentation prediction is obtained by averaging the denoised outputs from both branches, ensuring a robust and contextually informed result. Extensive experiments demonstrate that HopaDIFF achieves state-of-the-art performance on the proposed RHAS benchmark, significantly outperforming existing baselines in diverse evaluation settings.
@@ 2 Related Work
Referring Scene Understanding. Referring scene understanding, which leverages natural language to identify specific regions or objects within images or videos, has emerged as a valuable paradigm across various computer vision applications such as autonomous driving [55] and fine-grained action understanding [39]. The progress in this area has been greatly driven by the availability of high-quality, open-source datasets and benchmarks [5, 58, 49, 23, 45, 55, 7, 30], which provide standardized evaluation protocols and diverse referring scenarios. For instance, Li et al. [27] addressed referring image segmentation using a recurrent refinement network, whereas the CLEVR-Ref+ dataset introduced by Liu et al. [33] focused on visual reasoning with structured referring expressions. More recently, Wu et al. [55] proposed a benchmark for referring multi-object tracking, further broadening the scope of this research direction. Referring video segmentation [32, 56, 21, 61, 10] is a related research direction, primarily targeting pixel-wise semantic segmentation of a specified object within trimmed videos. In contrast, our task focuses on per-frame action recognition to determine the temporal boundaries of different actions. Peng et al. [39] proposed the first benchmark to facilitate referring fine-grained human action recognition, yet their task is restricted to trimmed videos.
In this work, we for the first time, conduct research on referring human action segmentation, which aims at achieving concrete action understanding on untrimmed videos. Since existing human action segmentation datasets [3, 24, 47] focused on single-person scenarios and consist of action sequences recorded according to a restricted protocol, which are suitable for the exploration of referring human action segmentation, we collect the first RHAS dataset, i.e., RHAS133, from $133$ different movies in multi-person scenarios. Our contributed dataset will serve as an important research foundation for the RHAS task.
Video-based Human Action Segmentation. Human action segmentation is a challenging video understanding task that involves dividing long, untrimmed sequences into frame-level action segments [9, 37, 31]. Unlike action recognition on trimmed clips with single labels, segmentation requires precise frame-wise predictions to model temporal boundaries. Common approaches leverage temporal convolutional networks [25, 19, 28, 16, 29, 38, 60], graph neural networks [22, 59], and transformers [57, 52, 11, 12, 48, 36] to model long-range dependencies. Multi-stage reasoning [57, 1, 28, 51, 13, 38, 54] further refines predictions iteratively.
Recent methods such as FACT[37], sparse transformers[2], ASQuery [15], and diffusion models [31] have advanced the field. However, they are mostly designed for single-person settings with fixed action protocols, overlooking multi-person complexities.
In this work, we tackle referring human action segmentation by adapting recognition models with vision-language features (e.g., BLIP-2 [26], CLIP [43]) for textual guidance. We propose HopaDIFF, a diffusion-based model that fuses holistic and partial cues via cross-input gated attentional xLSTM and incorporates frequency-domain conditioning. Our method sets a new state of the art on the proposed RHAS133 dataset.
@@ 3 Benchmark
Dataset. We present RHAS133, the first dataset specifically curated for the task of referring human action segmentation. It comprises $133$ movies featuring diverse, multi-person scenarios. Annotations follow the AVA protocol [18], with the action label set extended from $80$ to $137$ fine-grained classes to capture a broader range of human behaviors. Textual references and action annotations are manually labeled and cross-validated by $6$ annotators with domain expertise to ensure high annotation quality. The dataset spans $33$ hours of video and includes $542$ annotated individuals. Notably, the textual references describe target individuals without disclosing their actions.
Table 1 compares RHAS133 with existing human action segmentation datasets. Prior datasets, such as SALADS50 [47], Assembly101 [44], EgoProceL [3], Breakfast [24], and GTEA [14], are typically limited to single-person settings, lack natural language references, and rely on predefined procedural action sequences. The per-action class statistics and word cloud of the textual reference annotation are shown in Fig. 2. In contrast, RHAS133 uniquely combines multi-person interactions, textual references, and fine-grained action annotations in unstructured, third-person video contexts. These properties make it a more comprehensive and challenging benchmark for advancing research in referring human action segmentation. Additional statistics are provided in the supplementary materials.
@@
SALADS50 [47] Ego-centric 4h 50 25
Assembly101 [44] Ego-centric 513h 101 53
EgoProceL [3] Ego-centric 62h 16 130
Breakfast [24] Third-person 77h 10 52
GTEA [14] Ego-centric 20h 20 4
@@
Baselines. To construct the first RHAS benchmark, we select baselines from both the human action segmentation and referring human action recognition domains to comprehensively evaluate performance under this novel setting. From the human action segmentation field, we include FACT [37], ActDiff [31], LTContent [2], and ASQuery [15], as they represent state-of-the-art methods employing diverse strategies such as dual-branch cross-attention (FACT), diffusion-based action segmentation model (ActDiff), sparse long-range transformer (LTContent), and query-based segmentation with boundary refinement (ASQuery). These models are chosen for their demonstrated ability to handle long, untrimmed videos and capture temporal dependencies, which are the key requirements for RHAS. In addition, we include RefAtomNet [39] from the referring human action recognition domain due to its effective use of vision-language models for textual grounding, which aligns well with the core challenge in RHAS. These selected baselines allow us to assess the capabilities and limitations of both traditional segmentation techniques and reference-guided approaches under the RHAS framework.
@@ 4 Methodology
Our proposed HopaDIFF is a diffusion-based action segmentation model, as shown in Fig. 3. Diffusion models [20, 46] aim to approximate the true data distribution $q$ through a parameterized model $p_{\theta}$. These models comprise two key components: the forward process, which progressively corrupts input data with Gaussian noise over $T$ time steps following a predefined variance schedule $\Gamma=\{\gamma_{t}\mid t\in\left[1,T\right]\}$, and the reverse process, which employs a neural network to iteratively reconstruct the data from noisy observations. During the forward process, the corrupted sample at time step $t$ is computed as $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}}$, where $\alpha_{t}=1-\gamma_{t}$ and $\bar{\alpha}_{t}$ denotes the cumulative product of $\alpha_{t}$ over time. This formulation allows $\mathbf{z}_{t}$ to be directly sampled from the original data $\mathbf{z}_{0}$ in closed form, enabling a reparameterized expression of $\mathbf{z}$ suitable for optimization. The reverse process seeks to reverse this diffusion by gradually denoising the noisy sample at $T$ step back to a clean sample at $t=0$ step. Each step of the reverse trajectory can be formulated as:
@@
Here, the timestep $t$ is sampled uniformly during training, and the loss encourages accurate reconstruction of $\mathrm{z}_{0}$ from $\mathrm{z}_{t}$. During inference, a sample $\mathbf{z}_{0}$ is generated by starting from Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and iteratively applying the denoising step. Iterating Eq. 2 from $t=T$ to $t=0$ yields a sample $\mathbf{z}_{0}$ from $p_{\theta}(\mathbf{z}_{0})$. This formulation serves as the foundation for the existing diffusion-based action segmentation approach [31] and our HopaDIFF, enabling flexible conditioning and iterative refinement in the temporal domain.

### v2

@@ Abstract
Action segmentation is a core challenge in high-level video understanding, aiming to partition untrimmed videos into segments and assign each a label from a predefined action set. Existing methods primarily address single-person activities with fixed action sequences, overlooking multi-person scenarios. In this work, we pioneer textual reference-guided human action segmentation in multi-person settings, where a textual description specifies the target person for segmentation. We introduce the first dataset for Referring Human Action Segmentation, i.e., RHAS133, built from $133$ movies and annotated with $137$ fine-grained actions with $33h$ video data, together with textual descriptions for this new task. Benchmarking existing action segmentation methods on RHAS133 using VLM-based feature extractors reveals limited performance and poor aggregation of visual cues for the target person. To address this, we propose a holistic-partial aware Fourier-conditioned diffusion framework, i.e., HopaDIFF, leveraging a novel cross-input gate attentional xLSTM to enhance holistic-partial long-range reasoning and a novel Fourier condition to introduce more fine-grained control to improve the action segmentation generation. HopaDIFF achieves state-of-the-art results on RHAS133 in diverse evaluation settings. The dataset and code are available at https://github.com/KPeng9510/HopaDIFF.
@@ 1 Introduction
Human action recognition I3D ; peng2024referring ; liu2022motion ; wang2022deformable ; peng2024navigating ; gao2024hypergraph is a core problem with applications in robotics peng2022transdarc , surveillance Surveillance , and healthcare peng2022should . While recent methods have shown strong performance, they typically rely on pre-trimmed video clips, which demand extensive manual pre-processing and limit applicability in real-time or continuous scenarios. Human action segmentation 2019_CVPR_Farha ; 2020_PAMI_Li ; 2020_ECCV_Wang ; 2021_BMVC_Yi , by contrast, operates on untrimmed video streams and aims to identify both action classes and their temporal boundaries. This setting is more challenging but aligns more closely with real-world use cases, where continuous and online interpretation of human behavior is essential.
@@
However, existing human action segmentation approaches lu2024fact ; liu2023diffusion ; bahrami2023much ; gan2024asquery are predominantly designed for single-person scenarios, where the performed actions follow a predefined sequence dictated by a specific task protocol, e.g., assembling a tool sener2022assembly101 based on fixed instructions or preparing a salad 50Salads according to a set menu. Such settings lack the complexity and variability present in real-world daily life scenarios, where multiple individuals may appear in the same untrimmed video, and the action sequences are not pre-defined but exhibit significant randomness and spontaneity. This limitation hinders the generalizability of current human action segmentation methods to more dynamic and unconstrained environments.
@@
In this work, we introduce a novel task, termed Referring Human Action Segmentation (RHAS), which, to the best of our knowledge, is the first to address action segmentation in multi-person untrimmed videos guided by textual references. We contribute the RHAS133 dataset, which comprises $133$ untrimmed videos featuring $542$ individuals, where each person of interest is annotated with a textual reference. This work is inspired by prior works on referring video and image understanding peng2024referring . These textual descriptions serve as guidance for the model to localize and segment the actions of the referred individuals, with fine-grained, multi-label action annotations provided at the frame level following the AVA dataset format gu2018ava , as shown in Fig. 1.
@@
We further adapt several well-established human action segmentation methods, i.e., FACT lu2024fact , ActDiff liu2023diffusion , ASQuery gan2024asquery , and LTContent bahrami2023much , as well as one referring human action recognition method, i.e., RefAtomNet peng2024referring , to the proposed RHAS setting and evaluate their performance on our newly constructed dataset, which serves as the first RHAS benchmark. However, the experimental results reveal that these baseline methods achieve limited performance under the RHAS framework. This can be attributed to the fact that existing human action segmentation approaches are primarily designed for single-person scenarios without textual guidance and are not equipped to handle unstructured action sequences without a fixed protocol in complex, multi-person movie environments. Meanwhile, RefAtomNet peng2024navigating , originally developed for referring action recognition in trimmed videos, lacks the capacity for temporal modeling and long video understanding required by the RHAS task.
@@
To improve referring human action segmentation, we introduce HopaDIFF, a novel framework that employs a holistic- and partial-aware Fourier conditioned diffusion strategy guided by untrimmed video and textual references. It consists of two branches: a holistic branch that captures global context from the full video, and a partial branch that leverages GroundingDINO liu2024grounding to detect the referred person, generating a cropped video stream for fine-grained representation using a Vision-Language Model (VLM), e.g., BLIP-2 li2023blip2 . To facilitate effective long video reasoning and holistic-partial aware information exchange, we introduce a novel module, i.e., HP-xLSTM, which fuses holistic and partial features using cross-input gate attentional xLSTM. Furthermore, to exert finer control over the generative process of the diffusion model for action segmentation, we incorporate an additional conditioning mechanism from the frequency domain based on the embeddings delivered by the HP-xLSTM. This is achieved by applying a Discrete Fourier Transform (DFT) along the temporal axis to extract frequency-based cues, which are combined with original feature conditions to guide the denoising process and are used to preserve temporal coherence and fine-grained details on motion dynamics by guiding the model to maintain consistency across both low and high-frequency components. The final segmentation prediction is obtained by averaging the denoised outputs from both branches, ensuring a robust and contextually informed result. Extensive experiments demonstrate that HopaDIFF achieves state-of-the-art performance on the proposed RHAS benchmark, significantly outperforming existing baselines in diverse evaluation settings.
To summarize, this paper makes three key contributions: (1) we introduce, for the first time, the task of Referring Human Action Segmentation (RHAS), paving the way for action segmentation in complex, multi-person scenarios guided by natural language reference; (2) we present RHAS133, the first dataset for this task, and establish a comprehensive benchmark using diverse evaluation settings and adapted baseline methods; (3) we propose HopaDIFF, a diffusion-based framework tailored for RHAS, featuring the novel HP-xLSTM module for long-term temporal modeling through cross-input gate attention between holistic and partial branches together with additional Fourier condition. Our method achieves state-of-the-art performance on the RHAS133 dataset.
@@ 2 Related Work
Referring Scene Understanding. Referring scene understanding, which leverages natural language to identify specific regions or objects within images or videos, has emerged as a valuable paradigm across various computer vision applications such as autonomous driving wu2023referring and fine-grained action understanding peng2024referring . The progress in this area has been greatly driven by the availability of high-quality, open-source datasets and benchmarks bu2022scene ; yu2016modeling ; vasudevan2018object ; khoreva2019video ; seo2020urvos ; wu2023referring ; dang2023instructdet ; lin2024echotrack , which provide standardized evaluation protocols and diverse referring scenarios. For instance, Li et al. li2018referring addressed referring image segmentation using a recurrent refinement network, whereas the CLEVR-Ref+ dataset introduced by Liu et al. liu2019clevr focused on visual reasoning with structured referring expressions. More recently, Wu et al. wu2023referring proposed a benchmark for referring multi-object tracking, further broadening the scope of this research direction. Referring video segmentation liu2024real ; wu2022language ; huang2025unleashing ; zhu2024exploring ; ding2022language ; fu2024objectrelator is a related research direction, primarily targeting pixel-wise semantic segmentation of a specified object within trimmed videos. In contrast, our task focuses on per-frame action recognition to determine the temporal boundaries of different actions. Peng et al. peng2024referring proposed the first benchmark to facilitate referring fine-grained human action recognition, yet their task is restricted to trimmed videos. In this work, we for the first time, conduct research on referring human action segmentation, which aims at achieving concrete action understanding on untrimmed videos. Since existing human action segmentation datasets egoprocel ; Breakfast ; 50Salads focused on single-person scenarios and consist of action sequences recorded according to a restricted protocol, which are suitable for the exploration of referring human action segmentation, we collect the first RHAS dataset, i.e., RHAS133, from $133$ different movies in multi-person scenarios with textual references. Our contributed dataset will serve as an important research foundation for the RHAS task.
Video-based Human Action Segmentation. Human action segmentation is a challenging video understanding task that involves dividing long, untrimmed sequences into frame-level action segments 2022_Arxiv_Ding ; lu2024fact ; liu2023diffusion ; fu2019embodied ; zhang2023importance . Unlike action recognition on trimmed clips with single labels, segmentation requires precise frame-wise predictions to model temporal boundaries. Common approaches leverage temporal convolutional networks 2017_CVPR_Lea ; 2020_3DV_Hampiholi ; 2020_PAMI_Li ; 2021_CVPR_Gao ; 2021_NC_Li ; 2022_PR_Park ; 2022_IJCNN_Zhang , graph neural networks 2020_CVPR_Huang ; 2022_Arxiv_Zhang , and transformers 2021_BMVC_Yi ; 2022_Arxiv_Wang ; 2022_Arxiv_Du ; 2022_NPL_Du ; 2022_MS_Tian ; 2023_Arxiv_Liu to model long-range dependencies. Multi-stage reasoning 2021_BMVC_Yi ; 2022_IVC_Aziere ; 2020_PAMI_Li ; 2020_NC_Wang ; 2019_CVPR_Farha ; 2022_PR_Park ; 2020_ECCV_Wang further refines predictions iteratively. Recent methods such as FACTlu2024fact , sparse transformersbahrami2023much , ASQuery gan2024asquery , and diffusion models liu2023diffusion have advanced the field. However, they are mostly designed for single-person settings with fixed action protocols, overlooking multi-person complexities. In this work, we tackle referring human action segmentation by adapting recognition models with vision-language features (e.g., BLIP-2 li2023blip2 , CLIP radford2021learning ) for textual guidance. We propose HopaDIFF, a diffusion-based model that fuses holistic and partial cues via cross-input gated attentional xLSTM and incorporates frequency-domain conditioning. Our method sets a new state of the art on the proposed RHAS133 dataset.
@@ 3 Benchmark
Dataset. We present RHAS133, the first dataset specifically curated for the task of referring human action segmentation. It comprises $133$ movies featuring diverse, multi-person scenarios. Annotations follow the AVA protocol gu2018ava , with the action label set extended from $80$ to $137$ fine-grained classes to capture a broader range of human behaviors. Textual references and action annotations are manually labeled and cross-validated by $6$ annotators with domain expertise to ensure high annotation quality. The dataset spans $33$ hours of video and includes $542$ annotated individuals. Notably, the textual references describe target individuals without disclosing their actions.
Table 1 compares RHAS133 with existing human action segmentation datasets. Prior datasets, such as SALADS50 50Salads , Assembly101 sener2022assembly101 , EgoProceL egoprocel , Breakfast Breakfast , and GTEA GTEA , are typically limited to single-person settings, lack natural language references, and rely on predefined procedural action sequences. The per-action class statistics and word cloud of the textual reference annotation are shown in Fig. 2. In contrast, RHAS133 uniquely combines multi-person interactions, textual references, and fine-grained action annotations in unstructured, third-person video contexts. These properties make it a more comprehensive and challenging benchmark for advancing research in referring human action segmentation. Additional statistics are provided in the supplementary materials.
@@
SALADS50 50Salads Ego-centric 4h 50 25
Assembly101 sener2022assembly101 Ego-centric 513h 101 53
EgoProceL egoprocel Ego-centric 62h 16 130
Breakfast Breakfast Third-person 77h 10 52
GTEA GTEA Ego-centric 20h 20 4
@@
Baselines. To construct the first RHAS benchmark, we select baselines from both the human action segmentation and referring human action recognition domains to comprehensively evaluate performance under this novel setting. From the human action segmentation field, we include FACT lu2024fact , ActDiff liu2023diffusion , LTContent bahrami2023much , and ASQuery gan2024asquery , as they represent state-of-the-art methods employing diverse strategies such as dual-branch cross-attention (FACT), diffusion-based action segmentation model (ActDiff), sparse long-range transformer (LTContent), and query-based segmentation with boundary refinement (ASQuery). These models are chosen for their demonstrated ability to handle long, untrimmed videos and capture temporal dependencies, which are the key requirements for RHAS. In addition, we include RefAtomNet peng2024referring from the referring human action recognition domain due to its effective use of vision-language models for textual grounding, which aligns well with the core challenge in RHAS. These selected baselines allow us to assess the capabilities and limitations of both traditional segmentation techniques and reference-guided approaches under the RHAS framework.
@@ 4 Methodology
Our proposed HopaDIFF is a diffusion-based action segmentation model, as shown in Fig. 3. Diffusion models DDPM ; DDIM aim to approximate the true data distribution $q$ through a parameterized model $p_{\theta}$. These models comprise two key components: the forward process, which progressively corrupts input data with Gaussian noise over $T$ time steps following a predefined variance schedule $\Gamma=\{\gamma_{t}\mid t\in\left[1,T\right]\}$, and the reverse process, which employs a neural network to iteratively reconstruct the data from noisy observations. During the forward process, the corrupted sample at time step $t$ is computed as $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\epsilon\sqrt{1-\bar{\alpha}_{t}}$, where $\alpha_{t}=1-\gamma_{t}$ and $\bar{\alpha}_{t}$ denotes the cumulative product of $\alpha_{t}$ over time. This formulation allows $\mathbf{z}_{t}$ to be directly sampled from the original data $\mathbf{z}_{0}$ in closed form, enabling a reparameterized expression of $\mathbf{z}$ suitable for optimization. The reverse process seeks to reverse this diffusion by gradually denoising the noisy sample at $T$ step back to a clean sample at $t=0$ step. Each step of the reverse trajectory can be formulated as:
@@
Here, the timestep $t$ is sampled uniformly during training, and the loss encourages accurate reconstruction of $\mathrm{z}_{0}$ from $\mathrm{z}_{t}$. During inference, a sample $\mathbf{z}_{0}$ is generated by starting from Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ and iteratively applying the denoising step. Iterating Eq. 2 from $t=T$ to $t=0$ yields a sample $\mathbf{z}_{0}$ from $p_{\theta}(\mathbf{z}_{0})$. This formulation serves as the foundation for the existing diffusion-based action segmentation approach liu2023diffusion and our HopaDIFF, enabling flexible conditioning and iterative refinement in the temporal domain.